# Textual Unlearning Gives a False Sense of Unlearning

### v2

@@ 4.3.1 Threat Model
In the white-box scenario, the adversary has access to the model weights both before ($\theta_{original}$) and after ($\theta_{unlearn}$) unlearning and aims to reconstruct the unlearned samples. Additionally, we assume the adversary knows the unlearning algorithm $U$ employed by the model developer. This assumption could be practical. First, the unlearning algorithm itself is not highly confidential, thereby posing a risk of disclosure (e.g., via a complicit developer). Second, model developers may be obligated to disclose details of the unlearning algorithm, performances, and other relevant information to collaborators to uphold transparency and ensure informed updates regarding model versions.
@@ 4.3.2 The Proposed Method
Suppose that an unlearned sample consists of text $x$ and label $y$, where $x$ is a sequence of $L$ tokens ($\langle tk_{1},\ldots,tk_{L}\rangle$) with each $tk_{l}$ representing a number in the vocabulary $\mathcal{V}$, and $y\in\mathbb{R}^{1\times C}$ corresponds to one of $C$ categories. Intuitively, the adversary randomly selects $L$ tokens from $\mathcal{V}$ and guesses a label $c$ among $C$ categories. The guessed sample is then evaluated by applying the unlearning algorithm $U$ to check whether it enables updating $\theta_{original}$ to $\theta_{unlearn}$. However, However, this approach is computationally infeasible: (1) The text is represented in a large discrete space, making the search challenging. The adversary faces the task of identifying the target unlearned sample from $C\times L^{|\mathcal{V}|}$ candidates, which is practically intractable. (2) The search process is poorly guided, as the weight differences before and after unlearning are not well incorporated. (3) The lack of constraints for effective search.
@@
  Inputs: original weights $\theta_{original}$, unlearned weights $\theta_{unlearn}$, unlearning algorithm $U$, token embedding dimension $\mathbb{R}^{1\times d}$, learning rate $\alpha$, regularization factor $\beta$.
  $\hat{x}\sim\mathbb{R}^{L\times d},\hat{y}\sim\mathbb{R}^{1\times C}\quad\text{candidate initialization}$
@@
where $x_{low}$ and $x_{up}$ are calculated in $W_{wordEbd}$. The detailed procedure is outlined in Algorithm  3.
@@ 5.1 Experimental Setups
Datasets: In this paper, we investigate unlearning mechanisms on fine-tuned LMs, since the fine-tuning datasets are more configurable and practical for evaluations (Maini et al., 2024). However, previous studies indicate that LMs might have “already seen” the fine-turned samples in public datasets during pre-training, potentially undermining the reliability of unlearning processes (Maini et al., 2024). To ensure rigorous evaluation, we manually construct two synthetic datasets, SynthPAI-age and SynthPAI-inc, enabling fine-tuning on previously “unseen” data. They are two binary text classification datasets for attribute inference tasks, where texts consist of comments generated by GPT-4 based agents during interactions, labels are commenters’ attributes, such as age (young or old) and income (low or high). A detailed introduction on datasets is presented in Appendix A. Besides, we focus on text classification task to ensure a fair evaluation of MIA-based auditing and attack methods (Pawelczyk et al., 2023), as MIA remains poorly defined and lacks robust metrics for text generation tasks (Duan et al., 2024).
@@ 5.2 Evaluations on Textual Unlearning Auditing
Figure 2: Auditing textual unlearning methods via U-LiRA+. The dashed lines and bars represent the auditing results on the original models and corresponding unlearned models, respectively.
In this subsection, we evaluate our proposed auditing method, U-LiRA+, across five unlearning methods under various settings, as shown in Figure 2. We employ the SOTA auditing method U-LiRA as a baseline (Hayes et al., 2024). Detailed implementation of the auditing methods is provided in Appendix D. We default the unlearning ratio to $1\%$. We employ the True-Positive Rate at a low False-Positive Rate ($TPR@0.1\%FPR$) as the audit metric, reflecting how confidently an adversary compromises individual privacy. For $\mathcal{M}_{unlearn}$, a lower value indicates that the adversary has less confidence in distinguishing audit samples (unlearned and unseen samples), implying more successful unlearning.
Through rigorous auditing of existing unlearning methods, our empirical results reveal that the current auditing method significantly overestimates the effectiveness of unlearning mechanisms. For instance, in the GA audit results on the OPT-1.3b and SynthPAI-age, the baseline suggests that GA substantially reduces TPR on the unlearned model (from $21.01\%$ to $1.63\%$). However, our U-LiRA+ method shows that TPR remains at $17.23\%$ on $\mathcal{M}_{unlearn}$, indicating that the adversary can still effectively distinguish audit samples. Moreover, while Retrain could achieve a thorough erasure, inexact unlearning methods typically present a false sense of unlearning. Across various settings, inexact unlearning methods fail to effectively erase audited samples compared to Retrain, as evidenced by the consistently high TPR observed with U-LiRA+.
@@ 5.3 Exploring Black-Box Privacy Risks against Textual Unlearning Systems
5.3.1 TULA-MI in Strict Black-Box Cases
Table 1: TULA-MI in strict black-box cases. The metric is averaged $NTS@1NFS$ and the presented results are evaluated on the Pythia-1.4b model and SynthPAI-age dataset.
@@ 5.3.1 TULA-MI in Strict Black-Box Cases
We begin with the strict black-box scenarios where the adversary can only access output scores. To ensure a comprehensive evaluation, we examine three commonly used score functions: confidence ($f(x)_{y}$), cross-entropy loss ($-\log(f(x)_{y})$), and hinge loss ($\log(p/(1-p))$, $p=f(x)_{y}$). Additionally, we consider two unlearning ratios, $0.5\%$ and $1\%$ (the number of audited samples $N=32,64$ for SynthPAI-age and $24,48$ for SynthPAI-inc). Besides, we incorporate the metric $NTS@1NFS$, which quantifies the number of correctly inferred samples with only one mistake.
We present the results on Pythia-1.4b and SynthPAI-age dataset in Table 1, and results in other settings are provided in Appendix E. The results reveal that even with only access to output scores, the adversary can effectively infer membership information of the unlearned samples, with several samples correctly identified while just one incorrect. Moreover, Retrain are more vulnerable than inexact unlearning methods in such cases. As shown in Table 1, with a single incorrect guess, the adversary can infer more than 27 correct samples against the retrained model.
5.3.2 TULA-MI in Relaxed Black-Box Cases
Figure 3: TULA-MI in relaxed black-box cases.
We then move to the relaxed black-box cases, where the adversary can query the model to obtain logit vectors, is aware of the model architecture, and has access to an identically distributed auxiliary dataset. Under these conditions, we evaluate two baseline methods: LiRA (Carlini et al., 2022) and ULeaks (Chen et al., 2021). Detailed descriptions and implementations of these methods are provided in Appendix E. With a wrong guess below $0.1\%$, TULA-MI accurately infers membership for over $18\%$ of the samples, highlighting significant privacy risks. Compared to ULeaks, our method achieves superior performance by employing a more accurate per-sample attack strategy. While LiRA also utilizes a per-sample attack and fits normal distributions for unlearned and unseen outputs, its effectiveness is hindered by redundancy and overlap among textual samples, making it challenging to fit independent distributions for reliable discrimination. In contrast, our method explicitly learns complex decision boundaries for target samples in unlearned and unseen cases, enabling more robust MIA.
@@ 5.4 Exploring White-Box Privacy Risks against Textual Unlearning Systems
In this subsection, we examine the effectiveness of TULA-DR with white-box access to model weights before and after unlearning. To ensure a fair evaluation, we employ three metrics to assess reconstruction performance: ROUGE-1 (R-1), ROUGE-2 (R-2), and ROUGE-L (R-L) (Lin, 2004), which respectively measure the ratios of reconstructed unigrams, bigrams, and the longest-matching subsequence. Additionally, reconstructed examples are provided in Appendix I. We assume a default scenario where $\mathcal{M}_{original}$ is trained on $1000$ samples, with one sample subsequently unlearned. Furthermore, our method can be extended to batch unlearning scenarios, and the results are presented in Appendix G. For TULA-DR, the learning rate $\alpha$ is set to $0.1$ for the Pythia-1.4b model and $0.45$ for the OPT-1.3b model, and the regularization factor $\beta$ is fixed at $0.1$. Bseides, an ablation study is further provided in Appendix H.
As shown in Table 2, discrepancies in unlearning weights can result in substantial textual privacy leakage. Specifically, for the Pythia-1.4b model, over $80\%$ of the original words were accurately reconstructed (R-1), while the reconstructed sentences preserved both semantic structure (R-2) and meaning (R-3). Besides, for the OPT-1.3b model, although the reconstructed sentences exhibited disordered semantic structure, an average of $40\%$ of the original words were still recovered, exposing much information.

### v3

@@ 4.3.1 Threat Model
In the white-box scenario, the adversary has access to the model weights both before ($\theta_{original}$) and after ($\theta_{unlearn}$) unlearning and aims to reconstruct the unlearned samples. Additionally, we assume the adversary knows the unlearning algorithm $U$ employed by the model developer. This assumption could be practical. First, the unlearning algorithm itself is not highly confidential, thereby posing a risk of disclosure (e.g., via a complicit developer). Second, model developers may be obligated to disclose details of the unlearning algorithm, performance, and other relevant information to collaborators to uphold transparency and ensure informed updates regarding model versions.
@@ 4.3.2 The Proposed Method
Suppose that an unlearned sample consists of text $x$ and label $y$, where $x$ is a sequence of $L$ tokens ($\langle tk_{1},\ldots,tk_{L}\rangle$) with each $tk_{l}$ representing a number in the vocabulary $\mathcal{V}$, and $y\in\mathbb{R}^{1\times C}$ corresponds to one of $C$ categories. Intuitively, the adversary randomly selects $L$ tokens from $\mathcal{V}$ and guesses a label $c$ among $C$ categories. The guessed sample is then evaluated by applying the unlearning algorithm $U$ to check whether it enables updating $\theta_{original}$ to $\theta_{unlearn}$. However, this approach is computationally infeasible: (1) The text is represented in a large discrete space, making the search challenging. The adversary faces the task of identifying the target unlearned sample from $C\times L^{|\mathcal{V}|}$ candidates, which is practically intractable. (2) The search process is poorly guided, as the weight differences before and after unlearning are not well incorporated. (3) The lack of constraints for effective search.
@@
  Inputs: original weights $\theta_{original}$, unlearned weights $\theta_{unlearn}$, unlearning algorithm $U$, token embedding dimension $\mathbb{R}^{1\times d}$, learning rate $\alpha$, regularization factor $\beta$, the maximum number of iterations $N$.
  $\hat{x}\sim\mathbb{R}^{L\times d},\hat{y}\sim\mathbb{R}^{1\times C}\quad\text{candidate random initialization}$
@@
where $x_{low}$ and $x_{up}$ are calculated in $W_{wordEbd}$. The detailed procedure is outlined in Algorithm  3. Besides, the Decoding()($\cdot$)( ⋅ ) function refers to transforming the reconstructed embeddings into text space. Specifically, we use the model’s embedding layer as a “vocabulary” and apply cosine similarity to match the indices of the embeddings and then convert them into readable texts via the tokenizer.
@@ 5.1 Experimental Setups
Datasets: In this paper, we investigate unlearning mechanisms on fine-tuned LMs, since the fine-tuning datasets are more configurable and practical for evaluations (Maini et al., 2024). We manually construct two synthetic datasets, SynthPAI-age and SynthPAI-inc, for the evaluations. They are two binary text classification datasets for attribute inference tasks, where texts consist of comments generated by GPT-4-based agents during interactions, labels are commenters’ attributes, such as age (young or old) and income (low or high). A detailed introduction on datasets is presented in Appendix A. Besides, we focus on the text classification tasks, because: (1) they enable us to design rigorous audit samples for rigorous unlearning auditing in the worst case. Specifically, the utilized mislabeled samples have been found to be most vulnerable to privacy leakage on image classification tasks (Aerni et al., 2024), and rigorous samples for text generation tasks are currently under-explored. (2) They ensure a fair evaluation on MIA-based auditing and attack methods (Pawelczyk et al., 2023), as MIA remains poorly defined and lacks robust metrics for text generation tasks (Duan et al., 2024).
Why synthetic datasets: Previous studies indicate that LMs might have “already seen” the fine-tuned samples in public datasets during pre-training, potentially undermining the reliability of unlearning processes (Maini et al., 2024). In addition, the real-world data can not meet our requirement, since: (1) pre-trained models are trained on vast amounts of real-world data, which is typically closed-sourced. (2) Recent studies indicate that we can not yet robustly detect whether a specific sample was used among massive pre-training data (Zhang et al., 2024a). Therefore, we utilize synthetic datasets to rigorously ensure the fine-tuning on previously “unseen” data. Specifically, the synthetic samples are entirely based on fictional scenarios, and generated by GPT-4-based agents released after the models we utilized. Through these efforts, we hope to ensure that the setups are rigorous and thus avoid being potentially misleading.
@@ 5.2 Evaluations on Textual Unlearning Auditing
Figure 3: Auditing unlearning methods on SynthPAI-inc dataset. The lines and bars represent the auditing results on the original models and corresponding unlearned models, respectively. OPT-1.3b model (Top), Pythia-1.4b model (Bottom).
In this subsection, we evaluate our proposed auditing method, U-LiRA+, across five unlearning methods on SynthPAI-inc dataset, as shown in Figure 3. We employ the SOTA auditing method U-LiRA as a baseline (Hayes et al., 2024). Detailed implementation of U-LiRA+ and results on SynthPAI-age dataset are provided in Appendix D. We default the unlearning ratio to $1\%$. We employ the True-Positive Rate at a low False-Positive Rate ($TPR@0.1\%FPR$) as the audit metric, reflecting how confidently an adversary compromises individual privacy. For $\mathcal{M}_{unlearn}$, a lower value indicates that the adversary has less confidence in distinguishing audit samples (unlearned and unseen samples), implying more successful unlearning.
Through rigorous auditing of existing unlearning methods, our empirical results reveal that the current auditing method significantly overestimates the effectiveness of unlearning mechanisms. For instance, in the audit results of KL method on the OPT-1.3b model in Figure 3, the baseline suggests that KL substantially reduces TPR on the unlearned model (from $12.25\%$ to $0.28\%$). However, our U-LiRA+ method shows that TPR remains at $41.60\%$ on $\mathcal{M}_{unlearn}$, indicating that the adversary can still effectively distinguish audit samples after unlearning. Moreover, while Retrain could achieve a thorough erasure, inexact unlearning methods typically present a false sense of unlearning. Across various settings, inexact unlearning methods fail to effectively erase audited samples compared to Retrain, as evidenced by the consistently high TPR observed with U-LiRA+.
@@ 5.3 Exploring Black-Box Privacy Risks against Textual Unlearning Systems
5.3.1 TULA-MI in Strict Black-Box Case
Table 1: TULA-MI in strict black-box cases. The metric is averaged $NTS@1FS$, and the presented results are evaluated on the Pythia-1.4b model and SynthPAI-age dataset.
@@ 5.3.1 TULA-MI in Strict Black-Box Case
[…]
We present the results on Pythia-1.4b and SynthPAI-age dataset in Table 1, and results in other settings are provided in Appendix E. The results reveal that even with only access to output scores, the adversary can effectively infer membership information of the unlearned samples, with several samples correctly identified while just one incorrect. Moreover, Retrain is more vulnerable than inexact unlearning methods in such cases. As shown in Table 1, with a single incorrect guess, the adversary can infer more than 27 correct samples against the retrained model.
5.3.2 TULA-MI in Relaxed Black-Box Case
Figure 4: MIAs in relaxed black-box case on SynthPAI-age dataset. OPT-1.3b model (Top), Pythia-1.4b model (Bottom).
@@ 5.3.2 TULA-MI in Relaxed Black-Box Case
We then move to the relaxed black-box cases, where the adversary can query the model to obtain logit vectors, is aware of the model architecture, and has access to an identically distributed auxiliary dataset. Under these conditions, we evaluate two baseline methods: LiRA (Carlini et al., 2022) and ULeaks (Chen et al., 2021). Detailed descriptions and implementations of these methods are provided in Appendix G.
As shown in Figure 4, we present the MIA results on SynthPAI-age dataset, and additional results on SynthPAI-inc dataset are included in Appendix F. With a wrong guess below $0.1\%$, TULA-MI accurately infers membership for over $18\%$ of the samples, highlighting significant privacy risks. Compared to ULeaks, our method achieves superior performance by employing a more accurate per-sample attack strategy. While LiRA also utilizes a per-sample attack and fits normal distributions for unlearned and unseen outputs, its effectiveness is hindered by redundancy and overlap among textual samples, making it challenging to fit independent distributions for reliable discrimination. In contrast, our method explicitly learns complex decision boundaries for target samples in unlearned and unseen cases, enabling more robust MIA.
In this subsection, we examine the effectiveness of TULA-DR with white-box access to model weights before and after unlearning. To ensure a fair evaluation, we employ three metrics to assess reconstruction performance: ROUGE-1 (R-1), ROUGE-2 (R-2), and ROUGE-L (R-L) (Lin, 2004), which respectively measure the ratios of reconstructed unigrams, bigrams, and the longest-matching subsequence. Additionally, reconstructed examples are provided in Appendix J. We assume a default scenario where $\mathcal{M}_{original}$ is trained on $1000$ samples, with one sample subsequently unlearned. Furthermore, our method can be extended to batch unlearning scenarios, and the results are presented in Appendix H. For TULA-DR, the learning rate $\alpha$ is set to $0.1$ for the Pythia-1.4b model and $0.45$ for the OPT-1.3b model, and the regularization factor $\beta$ is fixed at $0.1$. Besides, an ablation study is further provided in Appendix I.
As shown in Table 2, discrepancies in unlearning weights can result in substantial textual privacy leakage. Specifically, for the Pythia-1.4b model, over $80\%$ of the original words were accurately reconstructed, as the R-1 values show. Besides, the reconstructed sentences preserved both semantic structure and meaning with high R-2 and R-L values, respectively. Besides, for the OPT-1.3b model, although the reconstructed sentences exhibited disordered semantic structure, an average of $40\%$ of the original words were still recovered, exposing much information. Specifically, the adversary could still utilize the reconstructed words to infer the useful information of the original texts.